Fermi National Laboratory
 

Operation Summary for October 25, 2010 - October 27, 2010

 

Monday Oct. 25

The day shift began with Operations monitoring the Tevatron (TeV) store 8205 (96.7E30), the Recycler stash (259.7E10), and the Antiproton Source (Pbar) stack (28.2E10), with the Main Injector (MI) off due to a kicker (KPS5S) problem, with the Linac and Booster operating normally, with NuMI off for target work, with MiniBooNE taking beam, with the Meson FTBF CALICE T-978 experiment off due to work on a power supply (HP3US), and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, FESS personnel changed out a faulted transformer cabinet fuse for HP3US; Operations established beam to T-978 at 9 AM. Operations terminated store 8205 at 10 AM. Operators prepared D-Zero for an access while cryo system personnel repaired a wet engine flywheel at TeV sector B4. Operations put the TeV into HEP shot setup at 1:30 PM and established store 8208 at 2:30 PM with an average initial luminosity (AIL) of 333E30.

During the evening shift, Operations disabled beam to T-978 at 6:05 PM.

Tuesday Oct. 26

The midnight shift began with Operations monitoring store 8208 (103.94E30) and the stash (285.1E10), with Pbar stacking (27.1E10), and with MiniBooNE taking beam.

During the midnight shift, at 12:20 AM, operators found a Linac RF station (LRF1) modulator sparking; operators turned off the station and called in Linac personnel. Linac personnel found and replaced a cracked surge resistor inside the modulator. After the repairs, Operations had Linac online by 5:35 AM. Operations established beam to T-978 at 5:45 AM and then disabled their beam at 7:10 AM. Operations terminated store 8208 at 7:25 AM and put the TeV into shot setup at 7:35 AM.

During the day shift, Operations established store 8209 at 8:44 AM with an AIL of 329E30. Operations established beam to T-978.

During the evening shift, Operations disabled beam to T-978 at 6 PM. Operations terminated store 8209 at 10:35 PM and put the TeV into shot setup at 10:50 PM. On the first proton load, lifetime was poor; operators called in TeV personnel to investigate. Operations established store 8210 at 11:30 PM with an AIL of 307E30. Operators turned off the Pelletron after shot setup.

Wednesday Oct. 27

The midnight shift began with Operations monitoring store 8210 (238.97E30) and the stash (81E10), with Pbar stacking (17.5E10), and with MiniBooNE taking beam.

During the midnight shift, Operations established beam to T-978 at 5:12 AM. Operations disabled stacking at 5:50 AM for scheduled AP0 and PreVault accesses and work. Operations disabled the Booster and MiniBooNE beam at 6:05 AM. Operations turned off the Booster Gradient Magnet Power Supply for access at 7:06 AM. Later, RF personnel began working on Booster RF stations.

The Plans for Wednesday

The plans for today are to maintain the store. The Run Coordinator will allow Pbar to conduct some vacuum work that will take 10-12 hours due to the dismantling and then rebuilding of a block wall. Other machines will also conduct work during this time.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/28/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory